# What’s Left After Distillation? How Knowledge Transfer Impacts Fairness and Bias

### v1

@@ Abstract
Knowledge Distillation is a commonly used Deep Neural Network (DNN) compression method, which often maintains overall generalization performance. However, we show that even for balanced image classification datasets, such as CIFAR-100, Tiny ImageNet and ImageNet, as many as 41% of the classes are statistically significantly affected by distillation when comparing class-wise accuracy (i.e. class bias) between a teacher/distilled student or distilled student/non-distilled student model. Changes in class bias are not necessarily an undesirable outcome when considered outside of the context of a model’s usage. Using two common fairness metrics, Demographic Parity Difference (DPD) and Equalized Odds Difference (EOD) on models trained with the CelebA, Trifeature, and HateXplain datasets, our results suggest that increasing the distillation temperature improves the distilled student model’s fairness — for DPD, the distilled student even surpasses the fairness of the teacher model at high temperatures. This study highlights the uneven effects of distillation on certain classes and its potentially significant role in fairness, emphasizing that caution is warranted when using distilled models for sensitive application domains.
@@ 1 Introduction
  • Enhancements in fairness may not always align with improvements in the model’s generalization performance and a trade-off should be selected based on the specific application.
@@ 3 Related Work
Despite it’s prevalence, fewer studies have been conducted on understanding the impact on bias and fairness of distillation. Lukasik et al. (2021) showed that improvements in overall accuracy by self-distillation may come at the expense of harming accuracy on certain classes, and in a study by Chai et al. (2022), the influence of the weight $\alpha$ in the distillation loss (Equation 6) on model fairness was analyzed, revealing that an increase in $\alpha$ enhances fairness. In our research, we broaden the investigation into distillation, assessing the effects of increasing the temperature parameter on class-level performance and fairness in the student model.
@@ 5 Experimental Design and Results
The baseline results for teacher models are listed in Tables 1 and 5. The baseline non-distilled student models achieve lower test accuracies compared to their corresponding teacher models, as also listed in the tables, as is typically observed for smaller models empirically. Full experimental details can be found in Appendix C.
@@
Based on our earlier experiments, we understand that distillation influences the class bias. To quantify this change in bias and its impact on the model’s fairness, we examine additional datasets that possess demographic attributes, allowing us to measure their fairness metrics effectively. One of the datasets we use is CelebFaces Attributes dataset (CelebA) (Liu et al., 2015) which is a comprehensive face attributes dataset, comprising over 200K celebrity images, each annotated with 40 different attributes. For our analysis, we select one of these attributes to serve as a binary label, focusing on training our model to classify whether the celebrity is smiling (Extra experiments with eyeglasses label can be found in Appendix B, Table 8). The CelebA dataset is particularly suitable for this study as it includes two key attributes, ’Young’ and ’Male’, which we identify as protected demographic attributes. This selection makes CelebA an ideal choice for assessing the impact of distillation on fairness metrics in the context of these attributes.
@@
DS 8 92.66 $\pm$ 0.13 1.15 $\pm$ 0.05 3.29 $\pm$ 0.06 92.66 $\pm$ 0.12 5.22 $\pm$ 0.02 9.29 $\pm$ 0.04
DS 9 93.18 $\pm$ 0.08 1.08 $\pm$ 0.03 3.22 $\pm$ 0.04 93.18 $\pm$ 0.15 5.14 $\pm$ 0.04 9.22 $\pm$ 0.08
DS 10 92.77 $\pm$ 0.16 1.01 $\pm$ 0.04 3.10 $\pm$ 0.02 92.57 $\pm$ 0.11 4.98 $\pm$ 0.03 9.10 $\pm$ 0.04
Figure 4: Evaluation of Fairness Metrics for Distilled Students. Equalized Odds Difference (EOD)and Demographic Parity Difference (DPD)are reported in % and lower values indicate improved fairness. (a) illustrates fairness metrics for the CelebA dataset concerning the ’Young’ demographic attribute, and (b) metrics for the Trifeature dataset with regard to the ’color’ attribute. It is notable that the models are fairer for the Trifeature dataset compared to the CelebA dataset with lower values in metrics. The explanation lies in the fact that the Trifeature dataset maintains a balanced distribution of demographic attributes, while the CelebA dataset contains biases that mirror real-world disparities.
@@
We calculate the fairness metrics for baseline models and distilled student with different temperatures. The outcomes are presented in Figure 4, with full results in Appendix B, Figure 9, where an increase in temperature is correlated with improved fairness, having smaller demographic parity and equalized odds differences. The teacher baseline demonstrates superior fairness compared to the student baseline across both demographic attributes. It is notable that the fairness of the distilled student can surpass that of the teacher model as seen in the changes in the demographic parity difference across the Young attribute. This improvement suggests that the knowledge transferred from the teacher model aids the student model in paying greater attention to demographic attributes, thereby promoting more equitable treatment.
@@
We also apply our analysis to the Trifeature (Hermann and Lampinen, 2020) synthetic dataset, to further assess fairness metrics. This dataset comprises 224x224 RGB images, each featuring one of ten shapes, rendered in ten different textures and colored in ten hues. Example images from this dataset are shown in Appendix A, Figure 7. This dataset’s balanced attribute representation eliminates the skewness typically seen in real-world data, enabling a focused study on the direct effects of distillation on model fairness and bias. In this study, we classify ’shape’ as the label attribute, while ’color’ and ’texture’ serve as demographic attributes (Extra experiments with texture label can be found in Appendix B, Table 9). The explicit distinction among these attributes within the dataset offers a straightforward scenario for exploring the nuanced impact of distillation on fairness metrics. We utilize a ResNet-20 model as the teacher and a LeNet-5 (LeCun et al., 1998) model, adapted for color images, as the student. This training is conducted over a range of distillation temperatures. Additionally, a LeNet-5 model is trained from scratch as a non-distilled baseline for comparative analysis. Experimental details for CelebA and Trifeature can be found in Appendix C.
@@
In this dataset, where our target has 10 classes, we initially calculate fairness metrics separately for each class, treating the other classes collectively as a single ’other’ class, effectively creating a binary classification scenario for each analysis. This allows us to evaluate fairness metrics for each class against the aggregated rest, considering each of the demographic attributes. Subsequently, we aggregate these metrics to understand the overall fairness of the model by averaging the metrics across all classes. The pseudocode of our calculation is shown in Appendix C, Algorithm 1. The findings from this dataset are consistent with those from the CelebA dataset, demonstrating that an increase in the distillation temperature leads to enhancement in the student model’s fairness. Furthermore, the findings reaffirm that the fairness of the distilled student model can surpass that of the teacher baseline. Just as we observed for class bias, this trend does not hold at very high temperatures (T$=20,30,40$) with the teacher model having almost uniform softmax outputs (Appendix B, Table 9, and Table 8).
@@
Figure 5: Evaluation of Fairness Metrics for Distilled Students. Equalized Odds Difference (EOD)and Demographic Parity Difference (DPD)are reported in % and lower values indicate improved fairness. (a) illustrates fairness metrics for the HateXplain dataset concerning the ’gender’ demographic attribute, and (b) with regard to the ’race’ attribute.
@@ 6 Discussion
Knowledge Distillation is an increasingly popular method for training DNNs, maintaining generalization performance while reducing the model’s computational complexity However, our findings reveal that distillation can otherwise affect the distilled model — specifically we found that even for nominally balanced benchmark image datasets, including CIFAR-10/100, SVHN, Tiny ImageNet and ImageNet, as many as 10–41% and 10–32% of classes are statistically significantly affected when comparing class-wise accuracy between a teacher/distilled student and distilled student/non-distilled student model respectively, varying based on the dataset and temperature (Table 1, Table 5). Our results (Figure 3) strongly suggest that the relationship between temperature and the number of significantly affected classes is characterized by larger temperatures maintaining the class biases found in the teacher model, while lower temperatures result in class biases closer to that of the baseline non-distilled student model. As already well understood, based on the specific dataset and/or model, there may be a significant difference in generalization performance with temperature — our results (Figure 3) suggest bias also be considered in this trade-off. We stress that changes in class bias are not necessarily an undesirable outcome in themselves, without an understanding of how that change affects decisions made in the context of a model’s usage. Indeed our experimental results with two common fairness metrics, DPD and EOD, on three datasets with demographic groups: CelebA, Trifeature, and HateXplain suggest that by increasing the value of the temperature, a distilled student model’s fairness improves. For DPD, the distilled student even surpasses the fairness of the teacher model at high temperatures (Figure 4).
@@ Acknowledgments
We acknowledge the support of Alberta Innovates, the Natural Sciences and Engineering Research Council of Canada (NSERC), and Defence Research and Development Canada (DRDC). We are grateful for computational resources made available to us by Denvr Dataworks, Amazon, and the Digital Research Alliance of Canada. We also would like to acknowledge the helpful feedback of Sara Hooker and Mike Lasby.
@@ Appendix A Full Dataset Details
Algorithm 1 Fairness Metrics Calculation
Input: List of classes $C$, Demographic attribute $A$
$AggregateFairness\leftarrow 0$
for each class $c$ in $C$ do
     Create binary labels for class $c$:
     $AggregateFairness\leftarrow AggregateFairness+F_{c}$
end for
$\overline{F}\leftarrow\frac{AggregateFairness}{\text{len}(C)}$
return Overall Fairness Metric $\overline{F}$
@@ Appendix B Extra Experiments
Table 8: Fairness Metrics and Distillation on CelebA. The performance of teacher, Non-Distilled Student (NDS), and Distilled Student (DS) models with a range of temperatures $T$ on the CelebA dataset. Fairness metrics are presented with regard to the Young demographic attribute. Mean and std. dev. are over five random inits.
@@
Table 9: Fairness Metrics and Distillation on Trifeature dataset. The performance of teacher, Non-Distilled Student (NDS), and Distilled Student (DS) models with a range of temperatures $T$ on the Trifeature dataset. Fairness metrics are presented with regard to the shape and texture attributes. Mean and std. dev. are over five random inits.
Trifeature (Shape) Trifeature (Texture)
@@
Figure 9: Evaluation of fairness metrics for distilled students. (a) illustrates fairness metrics for the CelebA dataset with ’smiling’ label concerning the ’Male’ demographic attribute (Left) and the ’Young’ demographic attribute (Right). (b) presents fairness metrics for the Trifeature dataset with ’shape’ label with regard to the ’color’ attribute (Left) and the ’texture’ attribute (Right). Equalized odds and demographic parity differences are reported in % and lower values indicate improved fairness. It is notable that the models are fairer for the Trifeature dataset compared to the celebA dataset with lower values in metrics. The explanation lies in the fact that the Trifeature dataset maintains a balanced distribution of demographic attributes, while the CelebA dataset contains biases that mirror real-world disparities.

### v2

@@ Abstract
Knowledge Distillation is a commonly used Deep Neural Network (DNN) compression method, which often maintains overall generalization performance. However, we show that even for balanced image classification datasets, such as CIFAR-100, Tiny ImageNet and ImageNet, as many as 41% of the classes are statistically significantly affected by distillation when comparing class-wise accuracy (i.e. class bias) between a teacher/distilled student or distilled student/non-distilled student model. Changes in class bias are not necessarily an undesirable outcome when considered outside of the context of a model’s usage. Using two common fairness metrics, Demographic Parity Difference (DPD) and Equalized Odds Difference (EOD) on models trained with the CelebA, Trifeature, and HateXplain datasets, our results suggest that increasing the distillation temperature improves the distilled student model’s fairness, and the distilled student fairness can even surpass the fairness of the teacher model at high temperatures. Additionally, we examine individual fairness, ensuring similar instances receive similar predictions. Our results confirm that higher temperatures also improve the distilled student model’s individual fairness. This study highlights the uneven effects of distillation on certain classes and its potentially significant role in fairness, emphasizing that caution is warranted when using distilled models for sensitive application domains.
@@ 1 Introduction
  • Enhancements in fairness may not always align with improvements in the model’s generalization performance; a trade-off should be selected based on the specific application.
@@ 2 Background
In cases where the base rates of different demographic groups differ significantly, the Demographic Parity Difference (DPD) metric may not fully reflect fairness. Therefore, we recommend caution when interpreting DPD, and note that we also consider other fairness metrics, such as Equalized Odds Difference (EOD).
Individual Fairness
In addition, we extend our work to analyzing individual fairness, ensuring that similar individuals receive similar predictions. To evaluate this, we employ the Lipschitz condition proposed by Dwork et al. (2012), which quantifies fairness as the degree of variation in predictions with respect to input similarity:
$\mathcal{I}(f)=\mathbb{E}_{(x,x^{\prime})\sim P}\left[\frac{|f(x)-f(x^{\prime})|}{d(x,x^{\prime})}\right],$ (5)
where $d(x,x^{\prime})$ measures the similarity between inputs using cosine similarity in a pre-trained feature space. This metric captures whether a model provides consistent predictions for semantically similar inputs, ensuring fairness at an individual level. By incorporating this analysis, we present a more granular and interpretable view of fairness in knowledge distillation.
@@ 3 Related Work
Despite its prevalence, fewer studies have been conducted on understanding the impact on bias and fairness of distillation. Lukasik et al. (2021) showed that improvements in overall accuracy by self-distillation may come at the expense of harming accuracy on certain classes, and in a study by Chai et al. (2022), the influence of the weight $\alpha$ in the distillation loss (Equation 7) on model fairness was analyzed, revealing that an increase in $\alpha$ enhances fairness. In our research, we broaden the investigation into distillation, assessing the effects of increasing the temperature parameter on class-level performance and fairness in the student model.
@@ 4 Methodology
If a uniform drop occurs across all classes, the relative change in class-wise accuracy would remain consistent with the overall accuracy drop. In such a case, we would not reject the null hypothesis because the changes are not class-specific or biased, and there would be no evidence that distillation has a disproportionately larger effect on certain classes. However, if the accuracy drop is not uniform — i.e., some classes experience a larger drop in accuracy than others — this would indicate that distillation affects certain classes differently, and the relative difference between class-wise accuracy and overall accuracy would not be consistent. In such a case, we would reject the null hypothesis and conclude that the distillation process significantly impacts class-level accuracy.
@@ 5 Experimental Design and Results
The baseline results for teacher models are listed in Tables 1 and 6. The baseline non-distilled student models achieve lower test accuracies compared to their corresponding teacher models, as also listed in the tables, as is typically observed for smaller models empirically. Full experimental details can be found in Appendix C.
@@
Based on our earlier experiments, we understand that distillation influences the class bias. We caution that the experiments on CIFAR-100 and ImageNet do not directly concern fairness metrics since these datasets do not contain sensitive attributes. While temperature affects class-wise accuracy, the fairness metrics are not directly applicable in these cases, as these datasets are balanced and lack demographic attributes. To quantify this change in bias and its impact on the model’s fairness, we examine additional datasets that possess demographic attributes, allowing us to measure their fairness metrics effectively. One of the datasets we use is CelebFaces Attributes dataset (CelebA) (Liu et al., 2015) which is a comprehensive face attributes dataset, comprising over 200K celebrity images, each annotated with 40 different attributes. For our analysis, we select one of these attributes to serve as a binary label, focusing on training our model to classify whether the celebrity is smiling (Extra experiments with eyeglasses label can be found in Appendix B, Table 9). The CelebA dataset is particularly suitable for this study as it includes two key attributes, ’Young’ and ’Male’, which we identify as protected demographic attributes. This selection makes CelebA an ideal choice for assessing the impact of distillation on fairness metrics in the context of these attributes.
@@
DS 8 92.66 $\pm$ 0.13 1.15 $\pm$ 0.05 3.29 $\pm$ 0.06 92.66 $\pm$ 0.12 5.22 $\pm$ 0.02 9.05 $\pm$ 0.04
DS 9 93.18 $\pm$ 0.08 1.08 $\pm$ 0.03 3.22 $\pm$ 0.04 93.18 $\pm$ 0.15 5.14 $\pm$ 0.04 9.01 $\pm$ 0.08
DS 10 92.77 $\pm$ 0.16 1.01 $\pm$ 0.04 3.10 $\pm$ 0.02 92.57 $\pm$ 0.11 4.98 $\pm$ 0.03 8.86 $\pm$ 0.04
Figure 4: Evaluation of Fairness Metrics for Distilled Students in Computer Vision (CV). Equalized Odds Difference (EOD)and Demographic Parity Difference (DPD)are reported in % and lower values indicate improved fairness. (a) illustrates fairness metrics for the CelebA dataset with ’smiling’ label concerning the ’Young’ demographic attribute and (b) concerning the ’Male’ demographic attribute. (c) presents fairness metrics for the Trifeature dataset with ’shape’ label with regard to the ’color’ attribute and (d) with regard to the ’texture’ attribute. It is notable that the models are fairer for the Trifeature dataset compared to the CelebA dataset with lower values in metrics. The explanation lies in the fact that the Trifeature dataset maintains a balanced distribution of demographic attributes, while the CelebA dataset contains biases that mirror real-world disparities. As seen in the second column, the downward trend does not continue at very high temperatures (T$=20,30,40$), as the teacher model generates nearly uniform softmax outputs.
@@
We calculate the fairness metrics for baseline models and distilled student with different temperatures. We compare the performance across demographic groups against the non-distilled student model baseline. This baseline allows us to measure how distillation impacts the model’s fairness by showing whether it exacerbates or mitigates fairness issues that already existed in the baseline model. In some contexts, a small change in fairness metrics could be considered substantial, especially in sensitive domains (e.g., hiring algorithms, loan approvals, or medical diagnoses), where even slight improvements in fairness can have a significant social or ethical impact. On the other hand, in less critical applications, this change might not carry the same weight. Therefore, it’s important to interpret these changes based on the domain-specific impact. The outcomes are presented in Figure 4 and Table 2, where an increase in temperature is correlated with improved fairness, having smaller demographic parity and equalized odds differences. The teacher baseline demonstrates superior fairness compared to the student baseline across both demographic attributes. It is notable that the fairness of the distilled student can surpass that of the teacher model as seen in the changes in fairness metrics in Figure 4. This improvement suggests that the knowledge transferred from the teacher model aids the student model in paying greater attention to demographic attributes, thereby promoting more equitable treatment. This trend does not continue at very high temperatures (T$=20,30,40$), as the teacher model generates nearly uniform softmax outputs (Figure 4, and Appendix B Table 9).
Algorithm 1 Fairness Metrics Calculation for Trifeature Dataset
Input: List of classes $C$, Demographic attribute $A$
$AggregateFairness\leftarrow 0$
for each class $c$ in $C$ do
     Create binary labels for class $c$:
     $AggregateFairness\leftarrow AggregateFairness+F_{c}$
end for
$\overline{F}\leftarrow\frac{AggregateFairness}{\text{len}(C)}$
return Overall Fairness Metric $\overline{F}$
@@
We also apply our analysis to the Trifeature (Hermann and Lampinen, 2020) synthetic dataset, to further assess fairness metrics. This dataset comprises 224x224 RGB images, each featuring one of ten shapes, rendered in ten different textures and colored in ten hues. Example images from this dataset are shown in Appendix A, Figure 7. This dataset’s balanced attribute representation eliminates the skewness typically seen in real-world data, enabling a focused study on the direct effects of distillation on model fairness and bias. In this study, we classify ’shape’ as the label attribute, while ’color’ and ’texture’ serve as demographic attributes (Extra experiments with texture attribute can be found in Appendix B, Table 10). The explicit distinction among these attributes within the dataset offers a straightforward scenario for exploring the nuanced impact of distillation on fairness metrics. We utilize a ResNet-20 model as the teacher and a LeNet-5 (LeCun et al., 1998) model, adapted for color images, as the student. This training is conducted over a range of distillation temperatures. Additionally, a LeNet-5 model is trained from scratch as a non-distilled baseline for comparative analysis. Experimental details for CelebA and Trifeature can be found in Appendix C.
@@
In this dataset, where our target has 10 classes, we initially calculate fairness metrics separately for each class, treating the other classes collectively as a single ’other’ class, effectively creating a binary classification scenario for each analysis. This allows us to evaluate fairness metrics for each class against the aggregated rest, considering each of the demographic attributes. Subsequently, we aggregate these metrics to understand the overall fairness of the model by averaging the metrics across all classes. The pseudocode of our calculation is shown in Algorithm 1.
The findings from this dataset shown in Table 2 and Figure 4 are consistent with those from the CelebA dataset, demonstrating that an increase in the distillation temperature leads to enhancement in the student model’s fairness. Furthermore, the findings reaffirm that the fairness of the distilled student model can surpass that of the teacher baseline. Just as we observed for class bias, this trend does not hold at very high temperatures (T$=20,30,40$) with the teacher model having almost uniform softmax outputs (Figure 4, and Appendix B, Table 10).
@@
Figure 5: Evaluation of Fairness Metrics for Distilled Students in Natural Language Processing (NLP). Equalized Odds Difference (EOD)and Demographic Parity Difference (DPD)are reported in % and lower values indicate improved fairness. (a) illustrates fairness metrics for the HateXplain dataset concerning the ’gender’ demographic attribute, and (b) with regard to the ’race’ attribute. The teacher employed the BERT architecture, while the student used the DistilBERT architecture.
Table 4: Individual Fairness Metrics Across Datasets. Individual fairness scores for Teacher, Non-Distilled Student (NDS), and Distilled Student (DS) models across CelebA, Trifeature, and HateXplain datasets. Scores for DS models are reported for varying temperature values $T$.
Individual Fairness ↓
CelebA Trifeature HateXplain
Model Temp ResNet-50 / ResNet-18 ResNet-20 / LeNet-5 BERT-Base / DistilBERT
Teacher 0.0407 0.016 0.0320
NDS 0.124 0.0462 0.1078
DS 1 0.113 0.0422 0.0994
DS 2 0.104 0.0407 0.0985
DS 3 0.0908 0.0393 0.0927
DS 4 0.0906 0.0387 0.0882
DS 5 0.0886 0.0384 0.0823
DS 6 0.0799 0.0377 0.0768
DS 7 0.0753 0.0356 0.0727
DS 8 0.0712 0.0349 0.0689
DS 9 0.0701 0.0341 0.0681
DS 10 0.0697 0.0338 0.0654
@@
We also assess individual fairness across the CelebA, Trifeature, and HateXplain datasets, as shown in Table 4. The results indicate that as the temperature increases, the individual fairness of the distilled student model improves, with the best scores observed at higher temperature values. These findings suggest that distillation not only enhances fairness at the group level but can also improve individual fairness across various datasets. The individual fairness scores offer a more detailed view of how distillation affects the model’s behavior toward specific instances, highlighting the potential of distillation to reduce bias in both group and individual decisions.
It is notable that with very large values of the temperature (T=20,30,40) and $\alpha$ distillation hyper-parameters, the student model does not receive as much useful information from the teacher. As the temperature increases, the teacher’s soft labels become more uniform, reducing the distinctiveness of the information passed on to the student. This results in both a decrease in test accuracy and a negative impact on fairness metrics as seen in Table 9 and Table 10.
@@ 6 Discussion
Knowledge Distillation is an increasingly popular method for training DNNs, maintaining generalization performance while reducing the model’s computational complexity. However, our findings reveal that distillation can otherwise affect the distilled model — specifically we found that even for nominally balanced benchmark image datasets, including CIFAR-10/100, SVHN, Tiny ImageNet and ImageNet, as many as 10–41% and 10–32% of classes are statistically significantly affected when comparing class-wise accuracy between a teacher/distilled student and distilled student/non-distilled student model respectively, varying based on the dataset and temperature (Table 1, Table 6). Our results (Figure 3) strongly suggest that the relationship between temperature and the number of significantly affected classes is characterized by larger temperatures maintaining the class biases found in the teacher model, while lower temperatures result in class biases closer to that of the baseline non-distilled student model. As already well understood, based on the specific dataset and/or model, there may be a significant difference in generalization performance with temperature — our results (Figure 3) suggest bias also be considered in this trade-off. We stress that changes in class bias are not necessarily an undesirable outcome in themselves, without an understanding of how that change affects decisions made in the context of a model’s usage. Indeed our experimental results with two common fairness metrics, DPD and EOD, on three datasets with demographic groups: CelebA, Trifeature, and HateXplain suggest that by increasing the value of the temperature, a distilled student model’s fairness improves. The distilled student fairness can even surpass the fairness of the teacher model at high temperatures (Figure 4).
@@ Acknowledgments
We acknowledge the support of Alberta Innovates (ALLRP-577350-22, ALLRP-222301502), the Natural Sciences and Engineering Research Council of Canada (RGPIN-2022-03120, DGECR-2022-00358), and Defence Research and Development Canada (DGDND-2022-03120). This research was enabled in part by support provided by the Digital Research Alliance of Canada (alliancecan.ca). We are also grateful for computational resources made available to us by Denvr Dataworks and an Amazon Research Award. We also would like to acknowledge the helpful feedback of Sara Hooker, Mike Lasby, and Abbas Omidi.
@@ Appendix B Extra Experiments
Table 9: Fairness Metrics and Distillation on CelebA. The performance of teacher, Non-Distilled Student (NDS), and Distilled Student (DS) models with a range of temperatures $T$ on the CelebA dataset. Fairness metrics are presented with regard to the Young demographic attribute for binary Smiling classification and binary Eyeglasses classification. Mean and std. dev. are over five random inits.
@@
Table 10: Fairness Metrics and Distillation on Trifeature dataset. The performance of teacher, Non-Distilled Student (NDS), and Distilled Student (DS) models with a range of temperatures $T$ on the Trifeature dataset for shape classification. Fairness metrics are presented with regard to color attributes and texture attributes. Mean and std. dev. are over five random inits.
Trifeature (Color) Trifeature (Texture)